Evaluate dy/dx from the equation  when 3xy^2+4x/y-(y-5)/(2x^2-1)=11 when x=1


1
Expert's answer
2021-04-15T07:34:12-0400

"Given \ 3xy^2 +\frac{4x}{y}-\frac{(y-5)}{(2x^2-1)}=11...(1)\\
To \ find \ \frac{dy}{dx} , when \ x=1.\\
We \ use \ implicit \ differentiation \\
Substituting \ x=1 \ in \ (1), we \ get \\
3(1)y^2 +\frac{4(1)}{y}-\frac{(y-5)}{(2(1)^2-1)}=11\\
\Rightarrow 3y^2 +\frac{4}{y}-(y-5)=11\\
\Rightarrow \frac{3y^3 +4-y(y-5)}{y}=11\\
\Rightarrow 3y^3 +4-y(y-5)=11y \ (by\ cross \ multiplication)\\
\Rightarrow 3y^3 +4-y^2+5y-11y=0\\
\Rightarrow 3y^3 -y^2-6y+4=0\\" "Since \ the \ sum \ of \ the \ coefficients \ is \ equal \ to \ zero \\
y=1 \ is \ a \ root \ of \ 3y^3-y^2-6y+4=0\\
By \ synthetic \ division \ we \ get \ 3y^2 +2y-4 \ as \ a \ factor\\
\Rightarrow y=\frac{-1+\sqrt(13)}{3} \ and \ y=\frac{-1-\sqrt(13)}{3}\\
\therefore y=1,\ \frac{-1+\sqrt(13)}{3} \ and \ \frac{-1-\sqrt(13)}{3}\\""Differentiating (1), with \ respect \ to \ 'x',\ we \ get \\
\frac{d}{dx}(3x^2y +\frac{4x}{y}-\frac{(y-5)})=\frac{d}{dx}(11) \\
\Rightarrow 

\frac{d}{dx}(3x^2y)+\frac{d}{dx}(\frac{4x}{y})-\frac{d}{dx}(\frac{(y-5)}{(2x^2-1)})=\frac{d}{dx}(11)\\
\Rightarrow [(3x^2)\frac{d}{dx}(y) +3y\frac{d}{dx}(x^2)]+[\frac{y\frac{d}{dx}(4x)-4x\frac{d}{dx}(y)}{y^2}]\\
-[\frac{(2x^2-1)\frac{d}{dx}(y-5)-(y-5)\frac{d}{dx}(2x^2-1)}{(2x^2-1)^2}]=0\\
\Rightarrow [(3x^2)\frac{dy}{dx} +6xy]+[\frac{4y-4x\frac{dy}{dx}}{y^2}]\\
-[\frac{(2x^2-1)\frac{dy}{dx}-(y-5)(4x)}{(2x^2-1)^2}]=0\\
Regrouping \ the \ terms \ containing \ \frac{dy}{dx}, we get \\
[3x^2-\frac{4x}{y^2}-\frac{1}{(2x^2-1)}]\frac{dy}{dx}=[-6xy-\frac{4y}{y^2}-\frac{(y-5)(4x)}{(2x^2-1)^2}]\\
\Rightarrow \frac{dy}{dx}=\frac{[-6xy-\frac{4y}{y^2}-\frac{(y-5)(4x)}{(2x^2-1)^2}]}{[3x^2-\frac{4x}{y^2}-\frac{1}{(2x^2-1)}]}\\
Substituting \ x=1 \ and \ y=1 \ in \frac{dy}{dx}, we \ get \\
[\frac{dy}{dx}]_{x=1,y=1}=\frac{[-6(1)(1)-\frac{4(1)}{(1)^2}-\frac{(1-5)(4(1))}{(2(1)^2-1)^2}]}{[3(1)^2-\frac{4(1)}{(1)^2}-\frac{1}{(2(1)^2-1)}]}\\
=\frac{[-6-4+16]}{[3-4-1]}\\
=\frac{(6)}{(-2)}=-3\\
\therefore [\frac{dy}{dx}]_{x=1,y=1}=-3"